OpenStudy (insa):

a two-digit number is such that the sum of the digits is 11. when the number with the same digits reversed is subtracted from the number, the difference is 9. what is the number

OpenStudy (anonymous):

Let xy be the number then x+y =11 What other equations can you write between x and y?

OpenStudy (anonymous):

The original number is 10 x +y The reversed is 10y + x Can you finish it now?

OpenStudy (insa):

i dont get the seconed equation

OpenStudy (anonymous):

\[10 x +y -(10 y + x) =9\\ 10 x + y -10y -x =9\\ 9 x - 9y =9\\ x-y =1 \] Can you finish it now?

OpenStudy (anonymous):

\[ x+ y =11\\ x-y =1\\ 2 x = 12\\ x=6\\ y=5 \] The number is 65 6+5 =11 65 -56=9

OpenStudy (insa):

now i get that...thx :)

OpenStudy (anonymous):

np